# Reconstructing and resampling: a guide to utilising posterior samples from gravitational wave observationsA

### v1

@@ 2.4 Calibration
The strain has a known calibration uncertainty, which is represented in the frequency domain by a distribution on the phase and amplitude (see, e.g. Sun et al., 2020). For standard LVK analyses, Bilby marginalises over the calibration uncertainty by taking as input a calibration envelope from which it constructs a spline approximation with an associated set of calibration parameters. These are then sampled during inference to marginalise over the calibration uncertainty (Farr et al., 2014; Abbott et al., 2016b). The calibration envelope is packaged as part of the data release and can be read in and used to construct a prior on the spline nodes. A calibration model must also be provided to the interferometer instances that contain the strain data. A demonstration of how this is done is provided in the notebooks.
@@ 2.5 Demonstrations
In the data release, we provide notebooks demonstrating how to implement the setting described within this section for two analyses and calculate the reconstructed likelihood (both marginalised and non-marginalised) and prior. We also show how to extract the stored values of the likelihood and prior. We choose GW150914, the first observed binary black hole merger (Abbott et al., 2016a), and GW230627_015337, a high-SNR binary black hole observed during the first part of the fourth observing run (Abac et al., 2025e). For GW150914, we use the GWTC-2.1 analysis (Abbott et al., 2024; LIGO Scientific Collaboration and Virgo Collaboration, 2022), which, amongst other things, utilised upgraded waveform models relative to the initial studies. In Fig.˜1, we show histograms of the residual between the reconstructed and stored log-likelihood and log-prior; in Table˜1, we summarise fits to the residuals. In both instances, the residuals are centred on zero, but while for the prior reconstruction we can achieve floating-point precision, for the likelihood, the residuals values are scattered with random offsets below the level of one in a thousand. For the log-likelihood, we fit both a normal distribution and a Student’s T distribution, showing that the typical variations better fit the Student’s T distribution with heavy tails.
While we do not extend this analysis to all observed events, we believe the attached notebooks should allow users to reconstruct the likelihood and prior for any observed event where Bilby analyses have been produced by the LVK.
@@
GW230627_015337 Marginalised (0.555 142 630 616 956-0.555\,142\,630\,616\,956, 0.447 842 238 315 843 130.447\,842\,238\,315\,843\,13) (0.000 824 287 632 696 041 3-0.000\,824\,287\,632\,696\,041\,3, 0.069 039 381 079 734 310.069\,039\,381\,079\,734\,31) (0.262 303 901 294 686 060.262\,303\,901\,294\,686\,06, $1.509\,323\,070\,576\,332\,4\text{\times}{10}^{-12}$, $9.266\,943\,721\,383\,815\text{\times}{10}^{-11}$)
Not-marginalised (0.541 548 158 366 822 5-0.541\,548\,158\,366\,822\,5, 0.443 662 542 636 502 60.443\,662\,542\,636\,502\,6) (0.000 365 233 485 950 966 5-0.000\,365\,233\,485\,950\,966\,5, 0.089 301 837 075 391 670.089\,301\,837\,075\,391\,67) (0.259 005 424 619 900 170.259\,005\,424\,619\,900\,17, $1.986\,371\,475\,329\,743\,7\text{\times}{10}^{-06}$, 0.000 462 751 264 404 301 50.000\,462\,751\,264\,404\,301\,5)
@@ 3 Resampling the posterior
$w(\theta)=\frac{p(\theta)}{p^{\prime}(\theta)}\,.$ (6)
@@ 3.1 Rejection sampling
where $\mathrm{sup}_{\theta}f(\theta)$ is the supremum of $f(\theta)$ for all possible values of $\theta$; in the context of continuous functions with a maximum, the supremum is the maximum. RS starts by drawing a sample $\theta_{i}\sim p(\theta)$ and then defining an acceptance probability for the sample $p^{\prime}(\theta_{i})/Mp^{\prime}(\theta_{i})$ which is proportional to how well the proposed sample matches the target distribution $p^{\prime}(\theta)$ normalised by the envelope $Mp^{\prime}(\theta)$. Algorithmically, this is implemented by first drawing $u\sim\mathrm{Unif}(0,1)$ and then accepting the sample if
@@
The bound $M$ ensures that $p^{\prime}(\theta)\leq Mp(\theta)$ for all possible values of $\theta$ and is ideally chosen to be as small as possible while respecting Eq.˜8. Visually, in Fig.˜2, the closer the $Mg(x)$ curve is to the peak of $f(\theta)$, the more efficient the sampler, as the smaller the rejection region is. More formally, if we define RS efficiency as
@@
Hence, the smallest value of $M$ that satisfies Eq.˜8 maximises the efficiency. In practice, the optimal value of $M$ is unknown in advance and must be chosen instead. Typically, this is done by selecting
@@
which ensures the condition is met over the set of samples. However, it does not ensure the condition is met in general. (From Fig.˜2, it can be seen that if Eq.˜8 is not met, the posterior will be clipped for all $\theta$ where $p^{\prime}(\theta)>Mp(\theta)$)).
If $M$ is approximated by Eq.˜13, the acceptance criteria is
@@
This approximation break down for $\sigma\gtrsim 1$, but provides a useful estimate of the expected efficiency for $\sigma<1$. Furthermore, for $\sigma\ll 1$ we obtain
@@
In the straightforward approach described above, the number of samples produced by RS, $N^{\prime}$ is determined by the acceptance ratio (see Eq.˜10). However, we note that it is also possible to repeatedly apply the RS algorithm until a pre-defined number of samples is produced, $N_{t}$. If $N_{t}>N^{\prime}$, the resulting sample set will no longer be independent and identically distributed (IID), but this technique is nevertheless useful for smoothing histograms (however, see the discussion in the next section for alternatives).
@@ 3.2 Importance sampling
IS is an alternative Monte Carlo approach which, in its textbook definition, rather than creating a new set of equally-weighted samples (as done by RS) instead utilises the samples and associated weights directly. For example, to visualise the re-weighted posterior distribution, a histogram could be produced using the weights to modify each sample’s contribution to the histogram density. Or alternatively, for summary statistics, the weights can be used to modify the contribution from each sample. E.g., the weighted mean of a distribution can be computed as
@@
The ESS was introduced (Kish, 1965) as a measure of the loss of precision due to unequal weights. Specifically, it can be interpreted as the number of samples that would need to be drawn from the target distribution to match the variance of the estimate taken from the weighted distribution. However, the definitions given in Eq.˜20 are an approximation that can be computed from the samples themselves (see Kong (1992) for the derivation and Elvira et al. (2022) for a detailed discussion).
@@
and hence, considering Eq.˜20 and assuming the weights are log-normally distributed, the efficiency of IS can then be estimated from
@@
In Fig.˜3, we compare the predicted efficiency of the RS and IS approaches assuming the weights are log-normally distributed. This suggests that IS is more efficient for moderate values of $\sigma$, though both of course become highly inefficient as $\sigma$ approaches unity. However, we note that these efficiencies should not be taken at face value, as they present fundamentally different statistical concepts: while $\eta_{\rm RS}$ is the acceptance rate of the RS algorithm, $\eta_{\rm IS}$ measures the ratio of the effective sample size of the weighted sample set. Nevertheless, they help us to identify a key difference between RS and IS, namely a trade-off between smoothness and noise: while RS produces pure equally-weighted samples, the level of noise (given by the square root of the number of accepted samples) is larger compared to IS, which keeps the entire set of samples (maximising the smoothness of the resulting histograms).
Figure 3: The efficiency of the RS and IS approaches as predicted by Eq.˜16, Eq.˜17 and Eq.˜22.
A common issue with IS is the requirement to package the weights alongside the samples. Since many downstream methods assume a set of equally weighted samples, this can be problematic. An alternative is therefore to instead apply multinomial-IS in which we draw $N_{t}$ samples (with replacement) from the original samples $\{\theta_{i}\}$ with probabilities given by the normalised weights in Eq.˜19. $N_{t}$ could either be taken to be the size of the original sample set, or a more conservative choice is to take $N_{\rm ess}$. In either case, the resulting resampled posterior $\{\theta_{i}^{\prime}\}$ is no longer IID (because it will contain repeated samples), but it can present smoother histograms compared to RS. We note that multinomial-IS is identical in behaviour to repeated application of RS (as discussed at the end of the last Section). However, multinomial-IS is computationally more efficient.
@@ 3.3 Pareto-smoothing
When the distribution of weights has a heavy right tail, both RS and IS can be highly inefficient. This typically occurs because the proposal and target distribution differ. A technique commonly applied in the field of statistics (but not yet common in gravitational-wave astronomy, thourgh see Mould et al. (2025)) is the use of Pareto-smoothing, which modifies the right tail of the weight distribution to stabilise resampling and reduce the error of IS estimates (Vehtari et al., 2024). The approach was developed as an extension of IS and named Pareto-smoothed (PSIS). A generalised Pareto distribution is fitted to the logarithm of the largest weights (typically assigned by the user, with a typical value of the largest 20%). These weights are then replaced with the expected order statistics of the fitted distribution, and then the entire set of weights are re-normalised. This process smoothes the tails of the distribution, and enables the entire set of samples to be used in IS with improved efficiency. Pareto smoothing does induce a small bias, but this is often acceptable since only the tail of the distribution is affected. However, the algorithm yields a diagnostic $\hat{k}$, the tail index of the fitted distribution, that can be used to assess the impact. For $\hat{k}<0.5$, the bias is expected to be negligible and results can be trusted. For $0.5\leq\hat{k}<0.7$, care should be taken. And for $\hat{k}>0.7$, the estimates are unstable or have a high degree of bias. $\hat{k}$ can also be negative, indicating that the tail of the weight distribution decays faster than a Pareto distribution would, and smoothing will have a negligible impact.
@@ 3.4 Evidence estimation
Typically, alongside the initial sample set $\{\theta_{i}\}$, the Bayesian evidence, Eq.˜2, is also estimated during the stochastic sampling process. When resampling, we can use the weights, Eq.˜7, to obtain a Bayes factor between the primed and unprimed evidence values as follows. First, we write the definition of the primed evidence
@@
where we have used the definition of the generalised weight function, Eq.˜6. Now, we recognise that $\mathcal{L}(d|\theta,M)\pi(\theta|M)=\mathcal{Z}(d|M)p(\theta|d,M)$ and therefore
@@
Given a set of weights, Eq.˜27 can be used to perform a comparison of the evidence under the two differing assumptions. If an estimate of $\mathcal{Z}(d|M)$ is available, then multiplying by the Monte Carlo average provides a new estimate of $\mathcal{Z}^{\prime}(d^{\prime}|M^{\prime})$.
@@
For a given set of weights, computing $\sigma_{w}$ and inserting this into Eq.˜29 can be used to estimate the uncertainty on the Bayes factor calculated in Eq.˜27. If instead, the log-Bayes factor is calculated, Eq.˜28, then propagating the uncertainty:
@@
Starting from Eq.˜29, we can understand the performance in more detail by estimating $\sigma_{w}$ (the true variance of the weights) with the sample variance:
@@
where we have rearranged the expression in the second line. Next, recalling the definition of the ESS, we can rewrite Eq.˜20 as
@@
Now combining with Eq.˜32, we can identify that
@@
and hence the relative error in the Bayes factor, Eq.˜27, is
@@
This expression shows that when the efficiency is high, the variance in the estimator of the Bayes factor is small, and we recover the usual variance scaling $\propto 1/\sqrt{N}$. Meanwhile, when the efficiency is small, such that $N_{\rm ess}\ll N$, the variance instead scales as $\propto 1/\sqrt{N_{\rm ess}}$. In other words, not only is the $N_{\rm ess}$useful in determining the efficiency of IS, it also determines the accuracy of the Monte Carlo integration approach to estimating the Bayes factor, Eq.˜27.
@@ 4 Examples
In this Section, we will provide several simple examples of the use of resampling. We begin with a simple toy model that investigates the sampling approaches discussed in Section˜3, then move on to demonstrations from gravitational-wave astronomy.
@@ 4.1 Validating and exploring resampling methods
We concoct a scenario in which the posteriors are under-constrained $\sigma_{e}/\sigma_{p}=2$, but unbiased, with $\delta\mu_{e}=0$, and then generate many realisations of the posterior distribution. In Fig.˜5, we calculate a probability probabilty (PP) test (Sidery et al., 2014; Veitch et al., 2015; Romero-Shaw et al., 2020) showing that the concocted posterior distributions are scale-biased (do not follow the expected 1-to-1 relation) and fail a $p$-value test. We discuss the specifics of the PP test further in Appendix˜A and extend the investigation to also consider $\delta\mu_{e}\neq 0$ showing how bias and scale-bias can be diagnosed from the failure of the PP test.
@@
We then repeat the simulation study, varying $\sigma_{p}$ while keeping $\sigma_{e}$ fixed. In Fig.˜6, we plot the efficiencies as calculated directly from Eq.˜10 and Eq.˜21. This allows us to study the performance of the different methods in practice (though, as discussed in Section˜3.2, the efficiency measures of RS and IS are not an apples-to-apples comparison).
@@
Figure 6: Measures of the efficiency of four resampling algorithms applied to concocted posterior distributions, described in Section˜4.1. For RS and PSRS, we calculate the efficiency from the acceptance ratio of the resulting distribution as defined in Eq.˜10. For IS and PSIS, we calculate the efficiency as the ratio of the ESS to the number of original samples, as defined in Eq.˜21. On the right-hand axes (brown solid curve), we plot the Pareto-smoothing diagnostic $\hat{k}$. For all quantities, we repeat the simulation study multiple times and then plot the median and 90% interval across the simulated realisations. We provide a vertical dashed line at $\sigma_{e}/\sigma_{p}=2$, corresponding to the simulation studied in Fig.˜5. We also provide a horizontal dashed line at $\hat{k}=0.7$ on the right-hand axis, a threshold above which the Pareto smoothing diagnostics suggests the estimates may be unstable or have a high degree of bias.

### v2

@@ 2.4 Calibration
The strain has a known calibration uncertainty, which is represented in the frequency domain by a distribution on the phase and amplitude (see, e.g. Sun and others, 2020). For standard LVK analyses, Bilby marginalises over the calibration uncertainty by taking as input a calibration envelope from which it constructs a spline approximation with an associated set of calibration parameters. These are then sampled during inference to marginalise over the calibration uncertainty (Farr et al., 2014; Abbott and others, 2016b). The calibration envelope is packaged as part of the data release and can be read in and used to construct a prior on the spline nodes. A calibration model must also be provided to the interferometer instances that contain the strain data. A demonstration of how this is done is provided in the data release.
@@ 2.5 Demonstrations
In the data release, we provide notebooks demonstrating how to implement the setting described within this section for two analyses and calculate the reconstructed likelihood (both marginalised and non-marginalised) and prior. We also show how to extract the stored values of the likelihood and prior. We choose GW150914, the first observed binary black hole merger (Abbott and others, 2016a), and GW230627_015337, a high-SNR binary black hole observed during the first part of the fourth observing run (Abac and others, 2025e). For GW150914, we use the GWTC-2.1 analysis (Abbott and others, 2024; LIGO Scientific Collaboration and Virgo Collaboration, 2022), which, amongst other things, utilised upgraded waveform models relative to the initial studies. In Fig.˜1, we show histograms of the residual between the reconstructed and stored log-likelihood and log-prior; in Table˜2, we summarise fits to the residuals. In both instances, the residuals are centred on zero, but while for the prior reconstruction we can achieve floating-point precision, for the likelihood, the residual values are scattered with random offsets below the level of one in a thousand. For the log-likelihood, we fit both a normal distribution and a Student’s T distribution, showing that the typical variations better fit the Student’s T distribution with heavy tails.
While we do not extend this analysis to all observed events, we believe the notebooks in the data release should allow users to reconstruct the likelihood and prior for any observed event where Bilby analyses have been produced by the LVK.
@@
GW230627_015337 Marginalised (0.555 142 630 616 956-0.555\,142\,630\,616\,956, 0.447 842 238 315 843 130.447\,842\,238\,315\,843\,13) (0.000 824 287 632 696 041 3-0.000\,824\,287\,632\,696\,041\,3, 0.069 039 381 079 734 310.069\,039\,381\,079\,734\,31) ($1.509\,323\,070\,576\,332\,4\text{\times}{10}^{-12}$, 0.262 303 901 294 686 060.262\,303\,901\,294\,686\,06, $9.266\,943\,721\,383\,815\text{\times}{10}^{-11}$)
Not-marginalised (0.541 548 158 366 822 5-0.541\,548\,158\,366\,822\,5, 0.443 662 542 636 502 60.443\,662\,542\,636\,502\,6) (0.000 365 233 485 950 966 5-0.000\,365\,233\,485\,950\,966\,5, 0.089 301 837 075 391 670.089\,301\,837\,075\,391\,67) ($1.986\,371\,475\,329\,743\,7\text{\times}{10}^{-06}$, 0.259 005 424 619 900 170.259\,005\,424\,619\,900\,17, 0.000 462 751 264 404 301 50.000\,462\,751\,264\,404\,301\,5)
@@ 3 Resampling the posterior
$w(\theta)=\frac{p^{\prime}(\theta)}{p(\theta)}\,.$ (8)
@@ 3.1 Rejection sampling
where $\mathrm{sup}_{\theta}f(\theta)$ is the supremum of $f(\theta)$ for all possible values of $\theta$; in the context of continuous functions with a maximum, the supremum is the maximum. RS starts by drawing a sample $\theta_{i}\sim p(\theta)$ and then defining an acceptance probability for the sample $p^{\prime}(\theta_{i})/Mp(\theta_{i})$ which is proportional to how well the proposed sample matches the target distribution $p^{\prime}(\theta)$ normalised by the envelope $Mp^{\prime}(\theta)$. Algorithmically, this is implemented by first drawing $u\sim\mathrm{Unif}(0,1)$ and then accepting the sample if
@@
The bound $M$ ensures that $p^{\prime}(\theta)\leq Mp(\theta)$ for all possible values of $\theta$ and is ideally chosen to be as small as possible while respecting Eq.˜10. Visually, in Fig.˜2, the closer the $Mg(x)$ curve is to the peak of $f(\theta)$, the more efficient the sampler, as the smaller the rejection region is. More formally, if we define RS efficiency as
@@
Hence, the smallest value of $M$ that satisfies Eq.˜10 maximises the efficiency. In practice, the optimal value of $M$ is unknown in advance and must be chosen instead. Typically, this is done by selecting
@@
which ensures the condition is met over the set of samples. However, it does not ensure the condition is met in general. (From Fig.˜2, it can be seen that if Eq.˜10 is not met, the posterior will be clipped for all $\theta$ where $p^{\prime}(\theta)>Mp(\theta)$).
If $M$ is approximated by Eq.˜15, the acceptance criteria is
@@
This approximation breaks down for $\sigma\gtrsim 1$, but provides a useful estimate of the expected efficiency for $\sigma<1$. Furthermore, for $\sigma\ll 1$ we obtain
@@
In the straightforward approach described above, the number of samples produced by RS, $N^{\prime}$, is determined by the acceptance ratio (see Eq.˜12). However, we note that it is also possible to repeatedly apply the RS algorithm until a pre-defined number of samples is produced, $N_{t}$. If $N_{t}>N^{\prime}$, the resulting sample set will no longer be independent and identically distributed (IID), but this technique is nevertheless useful for smoothing histograms (however, see the discussion in the next section for alternatives).
@@ 3.2 Importance sampling
IS is an alternative Monte Carlo approach which, in its textbook definition, rather than creating a new set of equally-weighted samples (as done by RS), instead utilises the samples and associated weights directly. For example, to visualise the re-weighted posterior distribution, a histogram could be produced using the weights to modify each sample’s contribution to the histogram density. Or alternatively, for summary statistics, the weights can be used to modify the contribution from each sample. E.g., the weighted mean of a distribution can be computed as
@@
The ESS was introduced (Kish, 1965) as a measure of the loss of precision due to unequal weights. Specifically, it can be interpreted as the number of samples that would need to be drawn from the target distribution to match the variance of the estimate taken from the weighted distribution. However, the definitions given in Eq.˜22 are an approximation that can be computed from the samples themselves (see Kong (1992) for the derivation and Elvira et al. (2022) for a detailed discussion).
@@
and hence, considering Eq.˜22 and assuming the weights are log-normally distributed, the efficiency of IS can then be estimated from
@@
In Fig.˜3, we compare the predicted efficiency of the RS and IS approaches assuming the weights are log-normally distributed. We also add a comparison with a numerical example, validating the predictions. This suggests that IS is more efficient for moderate values of $\sigma$, though both of course become highly inefficient as $\sigma$ approaches unity. However, we note that these efficiencies should not be taken at face value, as they present fundamentally different statistical concepts: while $\eta_{\rm RS}$ is the acceptance rate of the RS algorithm, $\eta_{\rm IS}$ measures the ratio of the effective sample size of the weighted sample set. Nevertheless, they help us to identify a key difference between RS and IS, namely a trade-off between smoothness and noise: while RS produces pure equally-weighted samples, the level of noise (given by the square root of the number of accepted samples) is larger compared to IS, which keeps the entire set of samples (maximising the smoothness of the resulting histograms).
Figure 3: The efficiency of the RS and IS approaches as predicted by Eq.˜18, Eq.˜19 and Eq.˜24. To numerically validate the theoretical predictions, we add numerical results from a toy model where the target distribution is a standard normal $x\sim N(0,1)$, and the proposal is a shifted normal distribution $x^{\prime}\sim N(\mu,1)$. For this setup, the log-weight is $\mu x-\mu^{2}/2$, which has standard deviation $\sigma=|\mu|$ when sampling $x\sim N(0,1)$. By varying $\mu$, we can control $\sigma$ and compare empirical efficiency measurements against the theoretical predictions, demonstrating agreement. For further details, please see the software behind the figure in the data release (Ashton, 2026).
A common issue with IS is the requirement to package the weights alongside the samples. Since many downstream methods assume a set of equally weighted samples, this can be problematic. An alternative is therefore to instead apply multinomial-IS in which we draw $N_{t}$ samples (with replacement) from the original samples $\{\theta_{i}\}$ with probabilities given by the normalised weights in Eq.˜21. $N_{t}$ could either be taken to be the size of the original sample set, or a more conservative choice is to take $N_{\rm ess}$. In either case, the resulting resampled posterior $\{\theta_{i}^{\prime}\}$ is no longer IID (because it will contain repeated samples), but it can present smoother histograms compared to RS. We note that multinomial-IS is identical in behaviour to repeated application of RS (as discussed at the end of the last Section). However, multinomial-IS is computationally more efficient.
@@ 3.3 Pareto-smoothing
When the distribution of weights has a heavy right tail, both RS and IS can be highly inefficient. This typically occurs because the proposal and target distribution differ. A technique commonly applied in the field of statistics (but not yet common in gravitational-wave astronomy, thourgh see Mould et al. (2025)) is the use of Pareto-smoothing, which modifies the right tail of the weight distribution to stabilise resampling and reduce the error of IS estimates (Vehtari et al., 2024). The approach was developed as an extension of IS and named Pareto-smoothed (PSIS). A generalised Pareto distribution is fitted to the logarithm of the largest weights (typically assigned by the user, with a typical value of the largest 20%). These weights are then replaced with the expected order statistics of the fitted distribution, and then the entire set of weights is re-normalised. This process smoothes the tails of the distribution, and enables the entire set of samples to be used in IS with improved efficiency. Pareto smoothing does induce a small bias, but this is often acceptable since only the tail of the distribution is affected. However, the algorithm yields a diagnostic $\hat{k}$, the tail index of the fitted distribution, that can be used to assess the impact. For $\hat{k}<0.5$, the bias is expected to be negligible, and results can be trusted. For $0.5\leq\hat{k}<0.7$, care should be taken. And for $\hat{k}>0.7$, the estimates are unstable or have a high degree of bias. $\hat{k}$ can also be negative, indicating that the tail of the weight distribution decays faster than a Pareto distribution would, and smoothing will have a negligible impact.
@@ 3.4 Evidence estimation
Typically, alongside the initial sample set $\{\theta_{i}\}$, the Bayesian evidence, Eq.˜2, is also estimated during the stochastic sampling process. When resampling, we can use the weights, Eq.˜9, to obtain a Bayes factor between the primed and unprimed evidence values as follows. First, we write the definition of the primed evidence
@@
where we have used the definition of the generalised weight function, Eq.˜8. Now, we recognise that $\mathcal{L}(d|\theta,M)\pi(\theta|M)=\mathcal{Z}(d|M)p(\theta|d,M)$ and therefore
@@
Given a set of weights, Eq.˜29 can be used to perform a comparison of the evidence under the two differing assumptions. If an estimate of $\mathcal{Z}(d|M)$ is available, then multiplying by the Monte Carlo average provides a new estimate of $\mathcal{Z}^{\prime}(d^{\prime}|M^{\prime})$.
@@
For a given set of weights, computing $\sigma_{w}$ and inserting this into Eq.˜31 can be used to estimate the uncertainty on the Bayes factor calculated in Eq.˜29. If instead, the log-Bayes factor is calculated, Eq.˜30, then propagating the uncertainty:
@@
Starting from Eq.˜31, we can understand the performance in more detail by estimating $\sigma_{w}$ (the true variance of the weights) with the sample variance:
@@
where we have rearranged the expression in the second line. Next, recalling the definition of the ESS, we can rewrite Eq.˜22 as
@@
Now combining with Eq.˜34, we can identify that
@@
and hence the relative error in the Bayes factor, Eq.˜29, is
@@
This expression shows that when the efficiency is high, the variance in the estimator of the Bayes factor is small, and we recover the usual variance scaling $\propto 1/\sqrt{N}$. Meanwhile, when the efficiency is small, such that $N_{\rm ess}\ll N$, the variance instead scales as $\propto 1/\sqrt{N_{\rm ess}}$. In other words, not only is the $N_{\rm ess}$useful in determining the efficiency of IS, it also determines the accuracy of the Monte Carlo integration approach to estimating the Bayes factor, Eq.˜29.
@@ 4 Examples
In this Section, we will provide several simple examples of the use of resampling. We begin with a simple toy model that investigates the sampling approaches discussed in Section˜3, then move on to demonstrations from gravitational-wave astronomy. Details of how to reproduce these examples can be found in the data release (Ashton, 2026).
@@ 4.1 Validating and exploring resampling methods
We concoct a scenario in which the posteriors are under-constrained $\sigma_{e}/\sigma_{p}=2$, but unbiased, with $\delta\mu_{e}=0$, and then generate many realisations of the posterior distribution. In Fig.˜5, we calculate a probability probabilty (PP) test (Sidery and others, 2014; Veitch and others, 2015; Romero-Shaw and others, 2020) showing that the concocted posterior distributions are scale-biased (do not follow the expected 1-to-1 relation) and fail a $p$-value test. We discuss the specifics of the PP test further in Appendix˜A and extend the investigation to also consider $\delta\mu_{e}\neq 0$, showing how bias and scale-bias can be diagnosed from the failure of the PP test.
@@
We then repeat the simulation study, varying $\sigma_{p}$ while keeping $\sigma_{e}$ fixed. In Fig.˜6, we plot the efficiencies as calculated directly from Eq.˜12 and Eq.˜23. This allows us to study the performance of the different methods in practice (though, as discussed in Section˜3.2, the efficiency measures of RS and IS are not an apples-to-apples comparison).
@@
Figure 6: Measures of the efficiency of four resampling algorithms applied to concocted posterior distributions, described in Section˜4.1. For RS and PSRS, we calculate the efficiency from the acceptance ratio of the resulting distribution as defined in Eq.˜12. For IS and PSIS, we calculate the efficiency as the ratio of the ESS to the number of original samples, as defined in Eq.˜23. On the right-hand axes (brown solid curve), we plot the Pareto-smoothing diagnostic $\hat{k}$. For all quantities, we repeat the simulation study multiple times and then plot the median and 90% interval across the simulated realisations. We provide a vertical dashed line at $\sigma_{e}/\sigma_{p}=2$, corresponding to the simulation studied in Fig.˜5. We also provide a horizontal dashed line at $\hat{k}=0.7$ on the right-hand axis, a threshold above which the Pareto smoothing diagnostics suggest the estimates may be unstable or have a high degree of bias.